Q1       Suppose that investment (I) is $400 billion, private saving (S) is $400 billion, (autonomous) taxes (T) are $500 billion, exports (X) are $300 billion, and imports (M) are $200 billion.

(a)        What is the government expenditure on goods and services? (Hint: S=Yd-C)

(b)        What is the government budget balance?

            For parts (c) and (d), think of the loanable funds approach:

(c)        Is the government exerting a positive or negative impact on investment?

(d)        What fiscal policy action might increase investment and speed economic growth? Explain how the policy action would work.
Answer #1

1. a) From the circular flow of expenditure and income, we know that I = S+T- G+M-X. When rearranging it will be G = S- I+ T+M-X. Hence I is $400 billion, S is $400 billion,T is $500 billion, X is $300 billion, M is $200 billion.

G = 400- 400+500+200-300 = 400

Therefore,government expenditure on goods and services = $400 billion.

b) To find government budget balance, substract taxes from expenditure. It is T - G = 500 - 400 = 100

Hence, the government budget balance is $100 billion.

c) By increasing the supply of loanable funds, which lowers the real interest rate and increases the investment. Hence it is exerting a positive impact on investment.

d) A further increase in taxes or a decrease government expenditure helps in increasing the supply of loanable funds and thus investment increases. This increase in investment speed up the economic growth.